OpenStudy (anonymous):

what is the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6)

OpenStudy (superdavesuper):

general line equation: y=mx+c where m is the slope n c is the y-intercept if two lines are parallel, their slopes are the same. what is the slope of the line given y=-3x+12?

OpenStudy (anonymous):

-3x

OpenStudy (anonymous):

the slope of the given line is -3 when two lines are parallel their slopes are equal therefore the slope of the required line id also -3 now we know slope = m = -3 and point (-1, 6) use this formula to get the equation y-y1 = m(x-x1) here m = -3 , x1 = -1 and y1 = 6 substitute all these values in the formula and get your answer